# Puzzle Similarity: A Perceptually-guided No-Reference Metric for Artifact Detection in 3D Scene Reconstructions

### v1

@@ Results
Puzzle Similarity consistently achieves high correlation with human-perceived artifacts across all datasets, outperforming well-established full-reference (FR) metrics such as SSIM, L1, and L2 norms, even without access to direct ground-truth correspondences. Tables 1 & 2 report the average Pearson and Spearman correlations for each scene. Our dataset includes three images per scene, with artifact types varying per scene. For example, the garden scene has prominent black regions due to holes in the reconstruction, making artifact detection straightforward and leading to high correlation scores for most metrics. However, scenes like treehill, stump, and flowers exhibit artifacts in the form of blurry or unnatural textures while maintaining similar color distributions to the ground truth. In these more subtle cases, Puzzle Similarity significantly outperforms FR metrics. We include extended results and artifact examples in our Supplemental.
Table 1: Pearson Correlation between Image Metrics (full-reference and no-reference (NR)) and Human Perception per Dataset
Full-Reference L1 0.199 0.325 0.384 0.280 0.078 0.630 0.533 0.556 0.141 0.314 0.114 0.314
L2 0.163 0.293 0.232 0.166 0.085 0.546 0.396 0.444 0.123 0.220 0.081 0.223
SSIM [37] 0.188 0.325 0.370 0.342 0.251 0.603 0.596 0.536 0.231 0.468 0.343 0.490
F
LIP [2]
0.268 0.403 0.452 0.366 0.194 0.719 0.684 0.652 0.239 0.319 0.142 0.402
LPIPS (vgg)) [49] 0.112 0.190 0.119 0.275 0.216 0.307 0.212 0.183 0.073 0.264 0.111 0.230
LPIPS (alex)) [49] 0.059 0.122 0.383 0.103 0.134 0.528 0.247 0.177 0.102 0.322 0.221 0.189
LPIPS (squeeze)) [49] 0.108 0.289 0.256 0.263 0.169 0.211 0.304 0.395 0.158 0.132 0.153 0.139
FovVideoVDP [19] 0.370 0.457 0.561 0.470 0.310 0.659 0.826 0.752 0.370 0.584 0.339 0.594
NR CNNIQA [12] 0.075 0.167 0.239 0.068 0.158 0.408 0.322 0.345 0.366 0.400 0.131 0.014
PIQE [26] 0.265 0.266 0.254 0.116 0.586 0.443 0.490 0.151 0.527 0.202 0.375 0.100
PuzzleSim (ours) 0.584 0.527 0.567 0.483 0.590 0.637 0.745 0.662 0.460 0.597 0.702 0.555
Table 2: Spearman Correlation between Image Metrics (full-reference and no-reference (NR)) and Human Perception per Dataset.
Full-Reference L1 0.151 0.237 0.231 0.209 0.051 0.391 0.347 0.501 0.118 0.298 0.191 0.288
L2 0.155 0.249 0.234 0.218 0.054 0.402 0.353 0.513 0.122 0.296 0.192 0.292
SSIM [37] 0.160 0.120 0.252 0.413 0.244 0.540 0.273 0.496 0.279 0.468 0.402 0.435
F
LIP [2]
0.205 0.335 0.250 0.272 0.203 0.466 0.438 0.564 0.210 0.294 0.216 0.365
LPIPS (vgg)) [49] 0.093 0.097 0.081 0.281 0.207 0.155 0.138 0.159 0.051 0.236 0.099 0.171
LPIPS (alex)) [49] 0.027 0.093 0.169 0.091 0.106 0.238 0.168 0.094 0.048 0.240 0.233 0.132
LPIPS (squeeze)) [49] 0.057 0.180 0.158 0.233 0.137 0.065 0.162 0.250 0.167 0.101 0.133 0.064
FovVideoVDP [19] 0.342 0.312 0.454 0.422 0.295 0.534 0.646 0.735 0.382 0.555 0.364 0.475
NR CNNIQA [12] 0.085 0.205 0.221 0.163 0.131 0.254 0.373 0.379 0.316 0.397 0.221 0.095
PIQE [26] 0.208 0.409 0.292 0.182 0.459 0.229 0.620 0.195 0.376 0.225 0.277 0.173
PuzzleSim (ours) 0.468 0.392 0.382 0.501 0.428 0.428 0.658 0.603 0.306 0.541 0.548 0.441
4.3 Comparison with Visual Difference Predictors
The explicit model of low-level human vision in VDP models usually produces strong correlations between the metric and human assessment. We compare against the state-of-the-art FovVideoVDP [19]; while the metric can take into account higher-order perceptual cues like motion and eccentricity, we disable them as we are 1) analyzing single static frames and 2) not assuming specific viewing conditions such as display size or distance to the screen. Puzzle Similarity not only matches but often surpasses FovVideoVDP, particularly in texture-rich scenes like treehill, stump, and flowers, again while relinquishing direct references.
@@ 4.4 Comparison with No-Reference Metrics
Closer to our work, we can compare with other no-reference quality metrics capable of producing spatial quality maps [45, 12]. Puzzle Similarity demonstrates superior accuracy in artifact localization, as shown by the correlation values in Table 3. While PIQE performed well in certain scenes, its overall correlation with human perception was generally lower in others, reflecting a less robust alignment with perceived artifact localization. However, while our method relies on a small subset of images from a similar distribution to the target image (e.g., the training dataset on novel view synthesis of a specific scene), metrics like PIQE do not require any extra images and attempt to generalize to any input. Still, in the context of 3D reconstruction, our metric performs better than all tested no-reference metrics.
Table 3: Aggregated correlation between Image Metrics and Human Perception with mean and standard deviation across all datasets.
@@
Full-Reference L1 $0.322_{\pm 0.203}$ $0.251_{\pm 0.157}$
L2 $0.248_{\pm 0.175}$ $0.257_{\pm 0.158}$
SSIM [37] $0.395_{\pm 0.166}$ $0.340_{\pm 0.168}$
F
LIP [2]
$0.403_{\pm 0.216}$ $0.318_{\pm 0.150}$
LPIPS (vgg) [49] $0.191_{\pm 0.147}$ $0.147_{\pm 0.114}$
LPIPS (alex) [49] $0.216_{\pm 0.169}$ $0.137_{\pm 0.107}$
LPIPS (squeeze) [49] $0.215_{\pm 0.148}$ $0.142_{\pm 0.097}$
FovVideoVDP [19] $0.524_{\pm 0.191}$ $0.460_{\pm 0.165}$
NR CNNIQA [12] $0.224_{\pm 0.174}$ $0.237_{\pm 0.153}$
PIQE [26] $0.314_{\pm 0.188}$ $0.304_{\pm 0.167}$
PuzzleSim (ours) $0.592_{\pm 0.119}$ $0.475_{\pm 0.137}$
@@ 5 Application: Progressive Automatic Artifact Inpainting
Finally, we will showcase a possible application of our metric in automatic restoration of novel images from a reconstructed scene.
Whenever it is possible to establish a visual distribution (e.g. we have a training dataset available), we can recursively use our metric to automatically identify visual outliers in novel views and remove them through inpainting.
@@
To apply neural inpainting, we first need to create a binary mask from the quality map $\mathcal{Q}$, indicating the areas to be inpainted. This involves finding an optimal threshold $\tau$ that clearly distinguishes artifact regions. The effectiveness of inpainting depends on setting this mask carefully. If the mask is too large, the inpainting may inadvertently remove clean parts of the scene. If the mask is too small, artifacts might be left untouched. In order to automatically find a balanced threshold, we use a conservative, iterative approach to refine the test image based on the assumption that artifacts have below-average quality scores. For an initial threshold, we select $N=50$ candidate values, uniformly spaced between the lowest and mean quality scores that we use to threshold the quality map.
@@
with $i$, $h$, and $w$ representing indices where $i=0,\dots,N-1$, $h=1,\dots,H_{\mathcal{I}}$, and $w=1,\dots,W_{\mathcal{I}}$. We generate inpaintings using all $N$ masks and recompute PuzzleSim for each option. The quality of each inpainted candidate is evaluated by calculating the average quality difference within the inpainted region, denoted as $\delta_{i}$. To discourage overly large masks, we add a regularization term that penalizes them. Further details on the mathematical definitions of $\delta_{i}$ and the regularization term are provided in the Supplementary material.
We then select the candidate that maximizes $\delta$. After determining the initial threshold, we iteratively refine the inpainted image by drawing new thresholds close to the previous one. For each new threshold, we repeat the selection procedure, but with fewer candidates sampled within a narrow range around the previous threshold. The size of this interval depends on a hyperparameter $\alpha$ and the spread of quality scores. Keeping this range small ensures stable convergence and prevents excessive inpainting that could disrupt scene geometry.
If the upper limit of the interval is below the minimum quality value $\min_{h,w}\hat{\mathcal{Q}}^{(h,w)}$, we revert to the initial sampling method in Eq. 7 since an empty mask would be meaningless and cause division by zero when computing $\delta_{i}$.
Finally, we terminate the process if no further improvement is achieved (i.e., $\max_{i}\delta_{i}\leq 0$), returning the final inpainting result. This framework guarantees a monotonic improvement in PuzzleSim quality.
In Figure 5, we showcase several novel views from the reconstructed scenes garden, kitchen, and flowers using only a fraction of the original training views (20-30%). We process this artifact-ridden new view through the iterative inpainting framework presented above. Our method successfully detects and inpaints artifacts in the original reconstruction, producing high-quality inpainting consistent with the distribution of the original scenes.
@@ 6 Limitations and Future Work
While our method demonstrates promising results, there are several limitations to consider. Finding the maximum similarity with many other vectors becomes expensive as the number of reference images and image resolution rises. Performing approximate maximum search or fitting Gaussian mixture models in the embedding space can improve computational performance [48, 11]. Furthermore, our metric is empirically calibrated, but choosing the weights to combine layers and weighting in the channel dimension in a data-driven manner could advance the metric further. The resolution at which our metric can be utilized is currently limited by the CNN backbone’s generalizability to higher resolutions. Although it is differentiable, it is unlikely to produce valuable gradients due to the max operation across many vectors. The next step would be to explore softmax alternatives to make the metric more suitable for gradient-based optimization.
@@ 7 Conclusion
In this work, we have introduced Puzzle Similarity, a no-reference visual quality metric designed to detect and localize artifacts in novel views generated by 3D scene reconstruction methods. By leveraging learned patch statistics from input views, our method can generate spatial artifact maps without needing ground-truth references, which is a significant advantage for evaluating reconstructed scenes. Furthermore, we have provided a novel dataset of human-assessed quality and artifact detection specifically tailored for 3D scene reconstruction approaches.
Our evaluation demonstrates that Puzzle Similarity outperforms traditional full-reference metrics, such as SSIM and L2 norms, as well as spatial quality no-reference metrics, like PIQE, in capturing artifacts that align with human perception. Compared to sophisticated visual difference predictors like FovVideoVDP, which leverage explicit fitted models of low-level human vision, Puzzle Similarity achieves comparable or superior performance in complex texture-rich scenes, proving its robustness across a range of artifact types and scenarios, while relinquishing the necessity of direct references.
Additionally, We demonstrate our metric applied to the problem of automatic artifact inpainting, highlighting its potential for enhancing quality in scene reconstructions. Overall, Puzzle Similarity offers an effective solution for artifact localization in 3D reconstruction, paving the way for more perceptually aligned, reference-free quality assessment in computer vision and graphics, and we are hopeful it will power many more downstream applications in the future, such as few-shot reconstruction and guided acquisition.

### v2

@@ Results
Tab. 1 reports the average Pearson and Spearman correlations for each scene. Our dataset includes three images per scene, with varying artifact types per scene. For example, the garden scene has prominent black regions due to holes in the reconstruction, making artifact detection straightforward and leading to high correlation scores for most metrics. However, scenes like treehill, stump, and flowers exhibit artifacts in the form of blurry or unnatural textures while preserving similar color distributions to the ground truth. Puzzle Similarity consistently achieves high correlation with human-perceived artifacts across all datasets, retaining smaller variance across datasets than all comparable methods (See Tab. 2). The higher variance in the averaged results is mainly due to scenes having different types of artifacts, where some are harder to identify than others (e.g., it is easier to identify black areas than small diffuse blobs). Keeping a small variance implies robust performance across various artifact types. We include extended quantitative and qualitative results in our Supplemental.
@@ 4.3 Comparison with No-Reference Metrics
We compare with other No-Reference quality metrics capable of producing spatial quality maps [63, 14, 67, 58, 53]. CNNIQA [14] was applied on patches as described in their paper. We applied padding to avoid cropping the borders and bilinearly upsampled the final map. PIQE [33] already produces three different kinds of maps that we averaged. PAL4VST [67], PaQ-2-PiQ [58] produce maps and were applied as they described in their papers but bilinearly upsampled to match the human maps’ resolution.
Puzzle Similarity demonstrates superior accuracy in artifact localization, as shown by the correlation values in Tab. 1. PAL4VST and CNNIQA performed poorly, as expected, given their focus on detecting specific types of distortions that are not necessarily similar to reconstruction artifacts. While PIQE and PaQ-2-PiQ performed well in certain scenes, their overall correlation with human opinion was generally lower in others, reflecting a less robust alignment with human assessment. However, while our method relies on a small subset of images from a similar distribution to the target image (e.g., the training dataset on novel view synthesis of a specific scene), NR metrics do not require any extra images and attempt to generalize to any input.
4.4 Comparison with Cross-Reference Metrics
CrossScore is to our knowledge the only other CR metric besides us, also reliant on the training views. We also bilinearly upsampled its output to match human maps. We show comparisons to CrossScore in Tab. 1 and Tab. 2, in Fig. 4 and extended results in the Supplementary. We outperform CrossScore in most datasets and show better performance both on average and in terms of consistency (with a substantially smaller standard deviation among results). While the expensive domain-specific pretraining of CrossScore should, in theory, be superior to our feature-space patch matching leveraging general models pre-trained on collections of natural images, their reliance on SSIM as its target quality assessment metric limits its potential to accurately model human quality assessment, due to the well-known limitations of the metric in this regard [36, 37, 68]. Their DINOv2 encoder limits map resolution to 14x14 blocks, reducing artifact localization accuracy. Furthermore, our approach is markedly simpler, can leverage any CNN as feature encoder, and can be adapted to specific domains by simply changing this backbone, without re-training or distilling any component of the framework.
4.5 Comparison with Full-Reference Metrics
While unsuitable for the few-shot reconstruction and no-reference artifact identification tasks our work focuses on, we provide extensive comparisons to Full-Reference metrics as well, given that our dataset further provides reference images (which are not normally available for these tasks). On average we further outperform all FR metrics, while CrossScore falls behind FovVideoVDP. Thus, in the context of 3D reconstruction, our metric performs better than all tested NR, CR and FR metrics. Investigating why even the most advanced VDPs, rooted in complex models of low-level human visual processing, quantitatively fall behind us, remains a fascinating avenue for future work.
Table 2: Aggregated correlation between Image Metrics and Human Perception with mean and standard deviation across all datasets. Above the dashed line, we list NR, below CR metrics.
@@ 4.5 Comparison with Full-Reference Metrics
PAL4VST [67] $0.078_{\pm 0.112}$ $0.062_{\pm 0.085}$
CNNIQA [14] $0.144_{\pm 0.247}$ $0.130_{\pm 0.253}$
PIQE [33] $0.292_{\pm 0.222}$ $0.268_{\pm 0.221}$
PaQ-2-PiQ [58] $0.402_{\pm 0.178}$ $0.349_{\pm 0.225}$
[…]
PuzzleSim (ours) $0.615_{\pm 0.120}$ $0.474_{\pm 0.137}$
@@ 5 Application: Progressive Automatic Artifact Inpainting
Finally, we showcase an application of our metric in automatic restoration of novel views from a reconstructed scene. Whenever it is possible to establish a visual distribution (e.g. we have a training dataset available), we can recursively use our metric to automatically identify visual outliers in novel views and remove them through inpainting.
@@
To apply neural inpainting, we first need to create a binary mask from the quality map $\mathcal{Q}$, indicating the areas to be inpainted. This involves finding an optimal threshold $\tau$ that clearly distinguishes artifact regions. The effectiveness of inpainting depends on carefully setting this mask. If the mask is too large, the inpainting may inadvertently remove clean parts of the scene. If the mask is too small, artifacts might be left untouched. In order to automatically find a balanced threshold, we use a conservative, iterative approach to refine the test image based on the assumption that artifacts have below-average quality scores. For an initial threshold, we select $N=50$ candidate values, uniformly spaced between the lowest and mean quality scores that we use to threshold the quality map.
@@
with $i$, $h$, and $w$ representing indices where $i=0,\dots,N-1$, $h=1,\dots,H_{\mathcal{I}}$, and $w=1,\dots,W_{\mathcal{I}}$. With LaMa (big) [44] we generate inpaintings using all $N$ masks and recompute PuzzleSim for each option. The quality of each inpainted candidate is evaluated by calculating the average quality difference within the inpainted region, denoted as $\delta_{i}$. To discourage overly large masks, we add a regularization term that penalizes them. Further details on the mathematical definitions of $\delta_{i}$ and the regularization term are provided in the Supplementary material. We then select the candidate that maximizes $\delta$. After determining the initial threshold, we iteratively refine the inpainted image by drawing new thresholds close to the previous one. For each, we repeat the selection procedure, but with fewer candidates sampled within a narrow range around the previous threshold. The size of this interval depends on a hyperparameter $\alpha$ and the spread of quality scores. Keeping this range small ensures stable convergence and prevents excessive, disrupting inpainting. If the upper limit of the interval is below the minimum quality value $\min_{h,w}\hat{\mathcal{Q}}^{(h,w)}$, we revert to the initial sampling method in Eq. 7 as an empty mask would be meaningless and cause division by zero when computing $\delta_{i}$. Finally, we terminate the process if no further improvement is achieved (i.e., $\max_{i}\delta_{i}\leq 0$), returning the final inpainting result. This framework guarantees a monotonic improvement in PuzzleSim quality. In Fig. 5, we showcase several novel views from the reconstructed scenes garden, kitchen, and flowers using only a fraction of the original training views (20-30%). We process this artifact-ridden new view through the iterative inpainting framework presented above. Our method successfully detects and inpaints artifacts in the original reconstruction, producing high-quality inpainting consistent with the distribution of the original scenes.
@@ 6 Limitations and Future Work
While our method demonstrates promising results, there are some limitations to consider. Even with our optimized implementation, finding the maximum similarity for a great number of vectors becomes expensive as the number of reference images and image resolution rises (see Supplemental material for implementation and runtime analysis). Performing approximate maximum search or fitting Gaussian mixture models in the embedding space can improve computational performance [66, 13]. Furthermore, our metric is empirically calibrated, but choosing the weights to combine layers and weighting in the channel dimension in a data-driven manner could advance the metric further. The resolution at which our metric can be utilized is currently bound by the CNN backbone’s generalizability to higher resolutions. Currently, we can not support Vision Transformer backbones as it takes special care to maintain a limited receptive field. Although our metric is differentiable, it is unlikely to produce valuable gradients due to the max operation across many vectors. The next step would be to explore softmax alternatives to make the metric more suitable for gradient-based optimization.
@@ 7 Conclusion
We proposed Puzzle Similarity, a Cross-Reference visual quality metric for detecting and localizing artifacts in novel views of 3D scene reconstructions. By leveraging learned patch statistics from input views, our method generates spatial artifact maps without requiring ground-truth references, addressing a key challenge in evaluating reconstructed scenes. To enable the evaluation of Cross-Reference metrics, we also provide a dataset of human-assessed quality and artifact localization for 3D scene reconstruction.
Our evaluation shows that Puzzle Similarity outperforms all tested Full-Reference, Cross-Reference and No-Reference metrics in capturing artifacts aligned with human perception, demonstrating robustness across diverse artifact types and texture-rich scenes without requiring direct references. Furthermore, we apply our metric to automatic image restoration, illustrating its potential to enhance scene reconstruction quality. Puzzle Similarity provides an effective, perceptually aligned, reference-free solution for artifact localization, with promising applications in few-shot reconstruction and guided acquisition.